OpenStudy (anonymous):

Please help. 1. If 2x^2 + y^2 = −2 then evaluate (d^2y/dx^2) when x = 2 and y = 3. 2. Find (dy/dx) if f(x) = (x + 1)^2x. a. 2xln(x + 1) b. 2ln(x+1) + (2x/(x+1)) c. 2x(x + 1)(2x − 1) d. [2ln(x+1) + (2x/(x+1)](x+1)^2x Please check these answers: 3. Find y' if x = cos(y). For my answer I got: −sin^2(y) 4. Find the slope of the graph of the relation y^3 − xy = −6 at the point (7, 2). For my answer I got: 2/5

OpenStudy (anonymous):

@ganeshie8 @myininaya @dan815 @iambatman @e.mccormick @abb0t @Destinymasha @TheSmartOne @undeadknight26 @triciaal @inkyvoyd @Secret-Ninja @Joel_the_boss @mathmate @pooja195 @acxbox22 @linn99123 @helppleaase @hhelpplzzzz @Conqueror @bohotness @jagatuba @zpupster @Phebe @CausticSyndicalist @hugsnotughs @Here_to_Help15

OpenStudy (anonymous):

Any help would be much appreciated!

OpenStudy (anonymous):

@Lyrae @TheSmartOne @myininaya @CausticSyndicalist

OpenStudy (anonymous):

\[2x^2 + y^2 = −2\\ 4x+2yy'=0\]\[y'=-\frac{2x}{y}\]

OpenStudy (anonymous):

@satellite73 okay yes got that far then i do

OpenStudy (anonymous):